# Efficient Individually Rational Recommender System under Stochastic Order

### v1

@@ A.2 Binary Structure
Notice that $\mathcal{P}\cup\mathcal{P}^{\prime}$ includes $O(K^{2})$ actions, while $\textit{IR}(s)$ for a state $s$ is generally a convex polytope with infinitely many actions. Further, in every non-terminate state $s$, ${\textit{IR}(s)\cap(\mathcal{P}\cup\mathcal{P}^{\prime})\neq\emptyset}$. Next, we remind the reader of the definition of $\mathcal{P}$-valid policies.
Definition A.2 (Mirroring Definition 3.1).
An IR policy $\pi$ is $\mathcal{P}$-valid if for every non-terminal state $s\in\mathcal{S}$,
@@
Observe that $\mathcal{P}$ is a strict subset of all IR actions in the state $A$, which incorporate mixes of at most two arms. However, the set of IR actions $\textit{IR}(s)$ for $s\subseteq A$ may include distributions mixing several elements of $A$. Due to the convexity of $W(\pi,s)$ in $\pi(s)$ (see the elaborated representation of $W$ in Equation (7)), the GMDP exhibits a nice structural property, as captured by the following Proposition A.1.
Proposition A.1 (Mirroring Proposition 3.1).
@@
The proof of Proposition A.1 appears in Section H. Due to Proposition A.1, we shall focus on $\mathcal{P}$-valid policies. Such policies are easy to visualize using trees, as we exemplify next. [^7] In Figure 2.1 we illustrated the optimal policy using a graph that is not a tree. However, a tree structure serves better the presentation of our technical statements.
Example A.1.
We reconsider Example 2.1, but neglect the actual distributions (as we only care about the expected values). Let $A=\{a_{1},a_{2},a_{3},a_{4}\}$, with $\textnormal{pos}(A)=\{a_{1},a_{2}\}$ and $\textnormal{neg}(A)=\{a_{3},a_{4}\}$. Consider the tree description in Figure 2. The root of the tree is the set of all arms. At the root, the policy picks $\boldsymbol{p}_{1,3}$. The outgoing left edge represents the case the realized action is $a_{1}$, which happens w.p. $\boldsymbol{p}_{1,3}(a_{1})$. In such a case, the new state is $\{a_{2},a_{3},a_{4}\}$. With the remaining probability, $\boldsymbol{p}_{1,3}(a_{3})$, the new state will be $\{a_{1},a_{2},a_{4}\}$. Leaves of the tree are terminal states, where no further exploration could be done. For instance, in the leftmost leaf, $\{a_{3},a_{4}\}$, the only arms explored are $\{a_{1},a_{2}\}$. The two highlighted nodes represent the same state. Since the presented policy is $\mathcal{P}$-valid, it is stationary; hence, the policy acts exactly the same in these two nodes and their sub-trees.
@@
The policy exemplified in Figure 2 has an additional combinatorial property: In every state $s$, it takes an action according to some order of the arms. This property is manifested in the following Definitions A.3 and A.4.
Definition A.3 (Right-ordered policy).
@@
Definition A.4 (Left-ordered policy).
@@
Figure 2: The policy described in Example A.1. Every node represents a state (the mapping is onto, but not one-to-one). Outgoing left edges imply the coin flips resulted in an arm from $\textnormal{pos}(v)$, and outgoing right edges imply an arm from $\textnormal{neg}(v)$. Leaves correspond to terminal states, where no action could be taken.
A.3 Proof Overview
We are ready to prove Theorem 3.1. The main tool in our analysis is Lemma A.1. Lemma A.1 reveals a rather surprising feature of $Q$: $Q$ is policy independent.
Lemma A.1 (Mirroring Lemma 3.1).
@@ A.3 Proof Overview
The proof of Lemma A.1 appears in Section B. We stress that this lemma holds regardless of Assumption 2.1. Next, we leverage Lemma A.1 to prove the main technical result of the paper.
Theorem A.1.
Let $\pi^{\star}$ be a right-ordered, $\mathcal{P}$-valid policy with $\sigma^{\textnormal{right}}_{\pi^{\star}}$ ordered in decreasing expected value. Under Assumption 2.1, for every state $s\in\mathcal{S}=2^{A}$, it holds that $W(\pi^{\star},s)=W^{\star}(s)$.
In particular, Theorem A.1 implies that $W({\small\textnormal{{IREGB}}},s_{0})=W^{\star}(s_{0})$ and that Theorem 3.1 holds, since IREGB is right-ordered in decreasing expected value. The formal proof of Theorem A.1 is relegated to Section E.
Appendix B Proof of Lemma A.1
  • Proof of Lemma A.1.We prove the lemma by a two-dimensional induction on the number of arms in $\textnormal{pos}(s)$ and $\textnormal{neg}(s)$. We prove four base cases in Section D:
    • $\left|\textnormal{pos}(s)\right|=1$ and $\left|\textnormal{neg}(s)\right|\geq 2$ (Proposition D.1).
    • $\left|\textnormal{pos}(s)\right|\geq 2$ and $\left|\textnormal{neg}(s)\right|=1$ (Proposition D.2).
    • $\left|\textnormal{pos}(s)\right|\geq 2$ and $\left|\textnormal{neg}(s)\right|=2$ (Proposition D.3).
    • $\left|\textnormal{pos}(s)\right|=2$ and $\left|\textnormal{neg}(s)\right|\geq 2$ (Proposition D.4).
@@
Similarly, $Q^{*}_{{i^{\prime}},{j^{\prime}}}(U)=Q(\rho,U)$; hence, proving that $Q(\pi,U)=Q(\rho,U)$ entails Equality (10). Next, let $\textnormal{suffix}(\sigma^{\textnormal{left}}_{\pi})$ be the set of all non-empty suffices of $\sigma^{\textnormal{left}}_{\pi}$. Being left-ordered suggests that on-path [^9] These are terminal states that $\pi$ reaches to with positive probability. terminal states with all arms of $U_{<}$ explored of $\pi$ are of the form $(Z,\emptyset)$, where $Z\in\textnormal{suffix}(\sigma^{\textnormal{left}}_{\pi})$. Next, we factor $Q(\pi,U)$ recursively as follows: We factor $Q(\pi,U)$ into two terms, like in Equation (B). Following, for each term obtained, we ask whether the corresponding state excludes $\{a_{j^{*}},a_{j^{\prime}}\}$. If the answer is yes, we stop factorizing it, and move to the other terms. We do this recursively, until we cannot factor anymore, or we reached a terminal state. Using this factorizing process, we have  [^10] We stop factorizing if both $a_{j^{*}},a_{j^{\prime}}$ were observed; thus, $Z$ will never be the empty set.
@@
Next, we want to simplify the coefficients $\left(c^{\pi}_{Z}\right)_{Z}$. We remark that $c^{\pi}_{Z}$ is not simply the probability of reaching $(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$ from $s$, i.e., $\Pr(s\stackrel{{\scriptstyle\pi}}{{\rightsquigarrow}}(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\}))$. To clarify, consider a strict suffix $Z$, $1\leq\left|Z\right|<\left|\textnormal{pos}(A)\right|$, and the suffix $Z^{\prime}=Z\cup\{a_{l}\}$ for the minimal element $a_{l}\in U_{>}\setminus Z$ according to $\sigma^{\textnormal{left}}_{\pi}$,i.e., $a_{l}=\operatorname*{arg\,min}_{a\in U_{>}\setminus Z}\sigma^{\textnormal{left}}_{\pi}(a)$. In the factorization process that produced Equation (12), once we got the term $Q(\pi,Z^{\prime},U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$, we stopped factorizing any further; thus, $c^{\pi}_{Z}$ does not include the probability of reaching a node associated with $(Z^{\prime},U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$ and then following the left edge to $(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$. However, this probability is taken into account in $\Pr(s\stackrel{{\scriptstyle\pi}}{{\rightsquigarrow}}(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\}))$. Rather, $c^{\pi}_{Z}$ is the probability of reaching any node $v$ in the tree induced by $\pi$ with the following property: $v$ represents the state $(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$, while $a_{j^{\prime}}$ does not belong to the state represented by the parent of $v$. In the tree interpretation, $v$ should also be a right child of its parent (for instance, the left highlighted node in the tree in Figure 2). The following Proposition B.1 describes $\left(c^{\pi}_{Z}\right)_{Z}$ in terms of $Q$.
Proposition B.1.
@@
The proof of Proposition B.1 appears at the end of this proof. Notice that for every $Z$, $c^{\pi}_{Z}$ includes values of $Q$ with less arms than $U$ (besides, perhaps, the case where $\left|U_{>}\right|=2$ and $\left|Z\right|=1$ obtaining $Q(\pi,U_{>},\{a_{j^{*}},a_{j^{\prime}}\})$, but we cover this case in the bases cases); consequently, due to the inductive step
@@
where the last equality follows from mirroring Proposition B.1 for $(c^{\rho}_{Z})_{Z}$. Ultimately,
@@
We follow the same technique as in the previous step. Let $\pi$ be an ordered policy such that $\sigma^{\textnormal{left}}_{\pi}=(a_{i^{*}},a_{i^{\prime}},\dots)$, i.e., $\sigma^{\textnormal{left}}_{\pi}$ ranks $a_{i^{*}}$ first, $a_{i^{\prime}}$ second and then follows some arbitrary order on the remaining arms, and $\sigma^{\textnormal{right}}_{\pi}=(a_{j^{\prime}},\dots)$. In addition, we define the ordered policy $\rho$ with $\sigma^{\textnormal{left}}_{\rho}=(a_{i^{\prime}},a_{i^{*}},\dots)$, where the dots refer to any arbitrary order on the remaining elements of $U_{>}$, and $\sigma^{\textnormal{right}}_{\rho}=\sigma^{\textnormal{right}}_{\pi}=(a_{j^{\prime}},\dots)$. Using the inductive step and the same arguments as in Equation (B), it suffices to show that $Q(\pi,U)=Q(\rho,U)$. We factor $Q(\pi,U)$ recursively such that
@@
The proof of Proposition B.2 appears at the end of this proof. Notice that for every $Z$, $d^{\pi}_{Z}$ includes values of $Q$ with less arms than $U$ (besides, perhaps, the case where $\left|U_{<}\right|=2$ and $\left|Z\right|=1$ obtaining $Q(\pi,\{a_{i^{*}},a_{i^{\prime}}\},U_{<}\})$, but we cover this case in the bases cases); consequently, due to the inductive step
@@
where the last equality follows from mirroring Proposition B.2 for $(d^{\rho}_{Z})_{Z}$. Ultimately, by rearranging Equation (16) and invoking the inductive step, Equation (18) and the fact that $\sigma^{\textnormal{right}}_{\rho}=\sigma^{\textnormal{right}}_{\pi}$, we get
@@
which contradicts Equation (9); hence, the lemma holds. ∎
Appendix C Additional Statements for Lemma A.1
  • Proof of Proposition B.1.
    Figure 3: Illustration for Proposition B.1. The tree depicts $T(\pi)$. Nodes $v_{1}$ to $v_{6}$ are nodes whose sub-trees were pruned in the construction of $T$. Let $Z=U_{>}\setminus\{a_{i^{\prime}},a_{i_{1}}\}$ and $Z^{c}=\{a_{i^{\prime}},a_{i_{1}}\}$. The minimal element of $Z$, denoted $a_{i(Z)}$ in the proof, is $a_{i_{2}}$. The corresponding $c^{\pi}_{Z}$ is the probability to reach one of $\{v_{3},v_{5},v_{6}\}$, namely, $c^{\pi}_{Z}=\Pr(\{v_{3},v_{5},v_{6}\})$. In the tree $T$, we ignore sub-trees of nodes $v$ labeled with $\dots$” since these do not contribute to $c^{\pi}_{Z}$. Observe that the probability of reaching $v_{i}$, for $i\in\{1,\dots,6\}$ is the same in $T(\pi)$ and $T$. Finally, notice that $Q(\pi,U_{>}\setminus Z\cup\{a_{i(Z)}\},\{a_{j^{*}},a_{j^{\prime}}\})=\Pr(\{v_{1},v_{2},\dots,v_{6}\})$, and $Q(\pi,U_{>}\setminus Z,\{a_{j^{*}},a_{j^{\prime}}\})=\Pr(\{v_{1},v_{2},v_{4}\})$. Combining, we get that $c^{\pi}_{Z}=Q(\pi,U_{>}\setminus Z\cup\{a_{i(Z)}\},\{a_{j^{*}},a_{j^{\prime}}\})-Q(\pi,U_{>}\setminus Z,\{a_{j^{*}},a_{j^{\prime}}\})=\Pr(\{v_{3},v_{5},v_{6}\})$ as required.

### v2

@@ 0.A.2 Binary Structure
Notice that $\mathcal{P}\cup\mathcal{P}^{\prime}$ includes $O(K^{2})$ actions, while $\textit{MIR}(s)$ for a state $s$ is generally a convex polytope with infinitely many actions. Further, in every non-terminate state $s$, ${\textit{MIR}(s)\cap(\mathcal{P}\cup\mathcal{P}^{\prime})\neq\emptyset}$. Next, we remind the reader of the definition of $\mathcal{P}$-valid policies.
Definition 6 (Mirroring Definition 3)
A MIR policy $\pi$ is $\mathcal{P}$-valid if for every non-terminal state $s\in\mathcal{S}$,
@@
Observe that $\mathcal{P}$ is a strict subset of all MIR actions in the state $A$, which incorporate mixes of at most two arms. However, the set of MIR actions $\textit{MIR}(s)$ for $s\subseteq A$ may include distributions mixing several elements of $A$. Due to the convexity of $W(\pi,s)$ in $\pi(s)$ (see the elaborated representation of $W$ in Equation (7)), the GMDP exhibits a nice structural property, as captured by the following Proposition 5.
Proposition 5 (Mirroring Proposition 1)
@@
The proof of Proposition 5 appears in Section 0.H. Due to Proposition 5, we shall focus on $\mathcal{P}$-valid policies. Such policies are easy to visualize using trees, as we exemplify next. [^7] In Figure 1 we illustrated the optimal policy using a graph that is not a tree. However, a tree structure serves better the presentation of our technical statements.
Example 2
We reconsider Example 1, but neglect the actual distributions (as we only care about the expected values). Let $A=\{a_{1},a_{2},a_{3},a_{4}\}$, with $\textnormal{pos}(A)=\{a_{1},a_{2}\}$ and $\textnormal{neg}(A)=\{a_{3},a_{4}\}$. Consider the tree description in Figure 2. The root of the tree is the set of all arms. At the root, the policy picks $\boldsymbol{p}_{1,3}$. The outgoing left edge represents the case the realized action is $a_{1}$, which happens w.p. $\boldsymbol{p}_{1,3}(a_{1})$. In such a case, the new state is $\{a_{2},a_{3},a_{4}\}$. With the remaining probability, $\boldsymbol{p}_{1,3}(a_{3})$, the new state will be $\{a_{1},a_{2},a_{4}\}$. Leaves of the tree are terminal states, where no further exploration could be done. For instance, in the leftmost leaf, $\{a_{3},a_{4}\}$, the only arms explored are $\{a_{1},a_{2}\}$. The two highlighted nodes represent the same state. Since the presented policy is $\mathcal{P}$-valid, it is stationary; hence, the policy acts exactly the same in these two nodes and their sub-trees.
@@
The policy exemplified in Figure 2 has an additional combinatorial property: In every state $s$, it takes an action according to some order of the arms. This property is manifested in the following Definitions 7 and 8.
Definition 7 (Right-ordered policy)
@@
Definition 8 (Left-ordered policy)
@@
Figure 2: The policy described in Example 2. Every node represents a state (the mapping is onto, but not one-to-one). Outgoing left edges imply the coin flips resulted in an arm from $\textnormal{pos}(v)$, and outgoing right edges imply an arm from $\textnormal{neg}(v)$. Leaves correspond to terminal states, where no action could be taken.
0.A.3 Proof Overview
We are ready to prove Theorem 3.1. The main tool in our analysis is Lemma 2. Lemma 2 reveals a rather surprising feature of $Q$: $Q$ is policy independent.
Lemma 2 (Mirroring Lemma 1)
@@ 0.A.3 Proof Overview
The proof of Lemma 2 appears in Section 0.B. We stress that this lemma holds regardless of Assumption 1. Next, we leverage Lemma 2 to prove the main technical result of the paper.
Theorem 0.A.1
Let $\pi^{\star}$ be a right-ordered, $\mathcal{P}$-valid policy with $\sigma^{\textnormal{right}}_{\pi^{\star}}$ ordered in decreasing expected value. Under Assumption 1, for every state $s\in\mathcal{S}=2^{A}$, it holds that $W(\pi^{\star},s)=W^{\star}(s)$.
In particular, Theorem 0.A.1 implies that $W({\small\textnormal{{IREGB}}},s_{0})=W^{\star}(s_{0})$ and that Theorem 3.1 holds, since IREGB is right-ordered in decreasing expected value. The formal proof of Theorem 0.A.1 is relegated to Section 0.E.
Appendix 0.B Proof of Lemma 2
Proof (Proof of Lemma 2)
We prove the lemma by a two-dimensional induction on the number of arms in $\textnormal{pos}(s)$ and $\textnormal{neg}(s)$. We prove four base cases in Section 0.D:
  • $\left|\textnormal{pos}(s)\right|=1$ and $\left|\textnormal{neg}(s)\right|\geq 2$ (Proposition 8).
  • $\left|\textnormal{pos}(s)\right|\geq 2$ and $\left|\textnormal{neg}(s)\right|=1$ (Proposition 9).
  • $\left|\textnormal{pos}(s)\right|\geq 2$ and $\left|\textnormal{neg}(s)\right|=2$ (Proposition 10).
  • $\left|\textnormal{pos}(s)\right|=2$ and $\left|\textnormal{neg}(s)\right|\geq 2$ (Proposition 11).
@@ Appendix 0.B Proof of Lemma 2
Similarly, $Q^{*}_{{i^{\prime}},{j^{\prime}}}(U)=Q(\rho,U)$; hence, proving that $Q(\pi,U)=Q(\rho,U)$ entails Equality (10). Next, let $\textnormal{suffix}(\sigma^{\textnormal{left}}_{\pi})$ be the set of all non-empty suffices of $\sigma^{\textnormal{left}}_{\pi}$. Being left-ordered suggests that on-path [^9] These are terminal states that $\pi$ reaches to with positive probability. terminal states with all arms of $U_{<}$ explored of $\pi$ are of the form $(Z,\emptyset)$, where $Z\in\textnormal{suffix}(\sigma^{\textnormal{left}}_{\pi})$. Next, we factor $Q(\pi,U)$ recursively as follows: We factor $Q(\pi,U)$ into two terms, like in Equation (0.B). Following, for each term obtained, we ask whether the corresponding state excludes $\{a_{j^{*}},a_{j^{\prime}}\}$. If the answer is yes, we stop factorizing it, and move to the other terms. We do this recursively, until we cannot factor anymore, or we reached a terminal state. Using this factorizing process, we have  [^10] We stop factorizing if both $a_{j^{*}},a_{j^{\prime}}$ were observed; thus, $Z$ will never be the empty set.
@@
Next, we want to simplify the coefficients $\left(c^{\pi}_{Z}\right)_{Z}$. We remark that $c^{\pi}_{Z}$ is not simply the probability of reaching $(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$ from $s$, i.e., $\Pr(s\stackrel{{\scriptstyle\pi}}{{\rightsquigarrow}}(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\}))$. To clarify, consider a strict suffix $Z$, $1\leq\left|Z\right|<\left|\textnormal{pos}(A)\right|$, and the suffix $Z^{\prime}=Z\cup\{a_{l}\}$ for the minimal element $a_{l}\in U_{>}\setminus Z$ according to $\sigma^{\textnormal{left}}_{\pi}$,i.e., $a_{l}=\operatorname*{arg\,min}_{a\in U_{>}\setminus Z}\sigma^{\textnormal{left}}_{\pi}(a)$. In the factorization process that produced Equation (12), once we got the term $Q(\pi,Z^{\prime},U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$, we stopped factorizing any further; thus, $c^{\pi}_{Z}$ does not include the probability of reaching a node associated with $(Z^{\prime},U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$ and then following the left edge to $(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$. However, this probability is taken into account in $\Pr(s\stackrel{{\scriptstyle\pi}}{{\rightsquigarrow}}(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\}))$. Rather, $c^{\pi}_{Z}$ is the probability of reaching any node $v$ in the tree induced by $\pi$ with the following property: $v$ represents the state $(Z,U_{<}\setminus\{a_{j^{*}},a_{j^{\prime}}\})$, while $a_{j^{\prime}}$ does not belong to the state represented by the parent of $v$. In the tree interpretation, $v$ should also be a right child of its parent (for instance, the left highlighted node in the tree in Figure 2). The following Proposition 6 describes $\left(c^{\pi}_{Z}\right)_{Z}$ in terms of $Q$.
Proposition 6
@@
The proof of Proposition 6 appears at the end of this proof. Notice that for every $Z$, $c^{\pi}_{Z}$ includes values of $Q$ with less arms than $U$ (besides, perhaps, the case where $\left|U_{>}\right|=2$ and $\left|Z\right|=1$ obtaining $Q(\pi,U_{>},\{a_{j^{*}},a_{j^{\prime}}\})$, but we cover this case in the bases cases); consequently, due to the inductive step
@@
where the last equality follows from mirroring Proposition 6 for $(c^{\rho}_{Z})_{Z}$. Ultimately,
@@
We follow the same technique as in the previous step. Let $\pi$ be an ordered policy such that $\sigma^{\textnormal{left}}_{\pi}=(a_{i^{*}},a_{i^{\prime}},\dots)$, i.e., $\sigma^{\textnormal{left}}_{\pi}$ ranks $a_{i^{*}}$ first, $a_{i^{\prime}}$ second and then follows some arbitrary order on the remaining arms, and $\sigma^{\textnormal{right}}_{\pi}=(a_{j^{\prime}},\dots)$. In addition, we define the ordered policy $\rho$ with $\sigma^{\textnormal{left}}_{\rho}=(a_{i^{\prime}},a_{i^{*}},\dots)$, where the dots refer to any arbitrary order on the remaining elements of $U_{>}$, and $\sigma^{\textnormal{right}}_{\rho}=\sigma^{\textnormal{right}}_{\pi}=(a_{j^{\prime}},\dots)$. Using the inductive step and the same arguments as in Equation (0.B), it suffices to show that $Q(\pi,U)=Q(\rho,U)$. We factor $Q(\pi,U)$ recursively such that
@@
The proof of Proposition 7 appears at the end of this proof. Notice that for every $Z$, $d^{\pi}_{Z}$ includes values of $Q$ with less arms than $U$ (besides, perhaps, the case where $\left|U_{<}\right|=2$ and $\left|Z\right|=1$ obtaining $Q(\pi,\{a_{i^{*}},a_{i^{\prime}}\},U_{<}\})$, but we cover this case in the bases cases); consequently, due to the inductive step
@@
where the last equality follows from mirroring Proposition 7 for $(d^{\rho}_{Z})_{Z}$. Ultimately, by rearranging Equation (16) and invoking the inductive step, Equation (18) and the fact that $\sigma^{\textnormal{right}}_{\rho}=\sigma^{\textnormal{right}}_{\pi}$, we get
@@
which contradicts Equation (9); hence, the lemma holds.
Appendix 0.C Additional Statements for Lemma 2
Proof (Proof of Proposition 6)
Figure 3: Illustration for Proposition 6. The tree depicts $T(\pi)$. Nodes $v_{1}$ to $v_{6}$ are nodes whose sub-trees were pruned in the construction of $T$. Let $Z=U_{>}\setminus\{a_{i^{\prime}},a_{i_{1}}\}$ and $Z^{c}=\{a_{i^{\prime}},a_{i_{1}}\}$. The minimal element of $Z$, denoted $a_{i(Z)}$ in the proof, is $a_{i_{2}}$. The corresponding $c^{\pi}_{Z}$ is the probability to reach one of $\{v_{3},v_{5},v_{6}\}$, namely, $c^{\pi}_{Z}=\Pr(\{v_{3},v_{5},v_{6}\})$. In the tree $T$, we ignore sub-trees of nodes $v$ labeled with $\dots$” since these do not contribute to $c^{\pi}_{Z}$. Observe that the probability of reaching $v_{i}$, for $i\in\{1,\dots,6\}$ is the same in $T(\pi)$ and $T$. Finally, notice that $Q(\pi,U_{>}\setminus Z\cup\{a_{i(Z)}\},\{a_{j^{*}},a_{j^{\prime}}\})=\Pr(\{v_{1},v_{2},\dots,v_{6}\})$, and $Q(\pi,U_{>}\setminus Z,\{a_{j^{*}},a_{j^{\prime}}\})=\Pr(\{v_{1},v_{2},v_{4}\})$. Combining, we get that $c^{\pi}_{Z}=Q(\pi,U_{>}\setminus Z\cup\{a_{i(Z)}\},\{a_{j^{*}},a_{j^{\prime}}\})-Q(\pi,U_{>}\setminus Z,\{a_{j^{*}},a_{j^{\prime}}\})=\Pr(\{v_{3},v_{5},v_{6}\})$ as required.